OpenStudy (anonymous):

If ,B+3 over 7 = 2 over 5 then which of the following is true? 5b + 15 = 14 b+ 3 = 14 5b+ 3 = 14 b+ 15 = 14

OpenStudy (turingtest):

if ____ we are missing a word, no?

OpenStudy (anonymous):

hold on pkz

OpenStudy (anonymous):

hm, I think we're missing something here :)

OpenStudy (anonymous):

does that help

OpenStudy (turingtest):

does what help?

OpenStudy (anonymous):

it's the second one i think. because you need to get rid of the over 7 so you multiply each side by 7 and you're left with b+3=14

OpenStudy (turingtest):

oh you modified your post :P ic

OpenStudy (anonymous):

\[\frac{b+3}{7} = \frac{2}{5}\]

OpenStudy (anonymous):

Multiplying both sides by 7 gives \[b+3 = \frac{14}{5}\] You can take it from here =)

OpenStudy (anonymous):

(B+3) /7 = 2/ 5 5(B+3)= 7*2 5B+15=14

OpenStudy (anonymous):

i got it never mind

OpenStudy (anonymous):

OH 2/5 lol I thought it was just two

OpenStudy (anonymous):